OpenStudy (anonymous):

Simplify square root of 80 end root minus 3 square root of 63 end root square root of 125 minus square root of 175

OpenStudy (anonymous):

could you put it in math form

OpenStudy (anonymous):

\[\sqrt{80}-3\sqrt{63}-\sqrt{125}-\sqrt{175}\]

OpenStudy (anonymous):

combine the square root by subtraction

OpenStudy (anonymous):